# Awes, Laws, and Flaws From Today’s LLM Research

### v2

@@ 2.2 Did it Change Over Time?
Figure 2: Change over time for selected criteria. Between 2022 and 2024 there was a large (+19%) increase in the use of LLMs as evaluators, and declines on the presence of ethics disclaimers (-12%) and statistical tests (-4%). Non-English evaluation, the use of human evaluators, and limitations sections are steady.
@@ 2.3 Do Papers With Certain Criteria Get More Citations?
We split our corpus in two (texts with and without the criterion analysed), and then used a two-sample Kolmogorov-Smirnov test to determine the probability that both samples came from the same distribution. A high probability would imply that there is no significant difference between the samples, and we could conclude that the presence of the criterion does not impact the number of citations.
Concretely, the null hypothesis $H_{0}$ in this test is that both samples come from the same underlying distribution. Accepting (rather, being unable to reject) $H_{0}$ means that the distributions are statistically indistinguishable, and hence we are unable to conclude that likelihood of citation is impacted by the presence of the criterion. Otherwise rejecting $H_{0}$ means the distributions are then distinct, and hence we may conclude that the criterion does impact the number of citations received.
For our corpus we first calibrated the $p$-value to $<0.05$ for all criteria. This means we thresholded the statistic for two samples of lengths $m,n$ to the Kolmogorov-Smirnov test condition $D_{m,n}=\sqrt{-\ln{(p/2)}\cdot(\frac{n+m}{2mn})}$. If the test statistic (percentage citation difference) $D$ is $D>D_{m,n}$, we reject $H_{0}$; and we expect to be wrong about our conclusions 5% of the time.
We rejected $H_{0}$ in the presence of ethics (Figure 3) and limitations sections; the use of LLM and automatic evaluators; open-sourcing; and claims of reason capabilities. This means that the presence of these criteria on a paper had an impact on the citations it received. For other criteria (error breakdowns, evaluation of non-English languages, claims of emergence, and negative results) we were unable to reject $H_{0}$ and hence the presence of the criterion did not affect the citations received. Full results are in Table 8.
Figure 3: Results of a Kolmogorov-Smirnov test on papers with ethics sections (blue) and without (orange). The test statistic is $D=0.12$, larger than our calibrated $D_{m,n}=0.09$ at $p<0.05$. The $p$-value for the test is $p=0.016$. This indicates that $98.4$% of the samples contain a difference in citations of over $12\%$, which indicates that the samples are not from the same underlying distribution. Alternatively, if both samples were sampled from the same distribution, the probability that they are as far apart as observed ($D>12\%$) is 1.6%. Since they were sampled from the same distribution, we may conclude that the presence of ethics sections does impact the number of citations received.
@@ 2.4 Are We Getting Better or Worse?
Finally, we analysed the relationship between citations and the presence of our criteria as a yearly trend from 2021 onwards. We measure this relationship as the gap (absolute percentage difference) between citations of papers containing the criterion and these that do not, aggregated by year. We evaluate the change in this gap: positive changes to the gap imply that papers containing the criterion are cited more often. See Table 12 for examples and the full results.
We observed a positive change in the gap for ethics sections between 2022 and 2023 (+10%), but a noticeable drop between 2023 and 2024 (-50%; Figure 4). Other criteria had increases in 2022-2023 and drops in 2023-2024, such as presence of statistical tests (+41% versus -44%); limitations sections (+38% and -12%); and negative results (31% and -30%). Remark that 2024 gap drops are expected, given the recency of these papers.
What was unexpected were upticks: LLMs as evaluators (+12% to +29%) and non-English evaluations (-25% to 40%) both had remarkable increases in citations.
Figure 4: Yearly change in the citation percentage (gap) between papers with and without ethics sections. Between 2021 and 2023, papers with ethics sections were increasingly cited more often (e.g. gap was -26% in 2022 versus -16% in 2023), up to 2024. The change in the gap between 2023 and 2024 was -50%; though explainable by the recency of the papers surveyed.
@@ 3 Discussion
Holistically, we found that very few papers claiming SOTA addressed ethical considerations, and proportions in line with the literature in other fields for statistical tests and open-sourcing. Similarly, articles suggesting emergent behaviour were common, but coupling of these claims with statistical tests or error breakdowns were rare, also in line with arguments about the true capabilities of these models.
Further analysis demonstrated an inflationary increase in research (half of 2024 is double the volume of all of 2023), and worrisome yearly trends regarding the criteria present in the papers, such as fewer ethics disclaimers and open sourcing statements. Steady criteria, such as limitations sections, may be explained by it being a hard requirement in ACL conferences [40]. Others, such as the use of LLM evaluators, increased significantly, and most papers using this technology claimed SOTA results. Claims of reasoning were evaluated with LLMs and not humans; versus claims that they could not reason, typically done with human evaluation alone. That said, papers claiming SOTA results had fewer statistical tests to back their claims when compared to papers that did not claim SOTA; but high proportions of open-sourcing and versioning reports. The use of LLM evaluators could be explained by technological developments and propagation of experimental protocols: namely, GPT-4’s technical report was released in 2023. This model and its variants have become the standard for LLM-based evaluation regardless of the community’s consensus around their effectiveness, perhaps because of their ease of access. However, other research protocols (parameter declarations, versioning, and randomness), decreased slightly that same year, which suggests rushed research when using this technology.
The second half of our work focused on the relationship of citations and the presence of criteria. We observed a statistically significant difference in the citations received by a paper when presenting ethics, limitations, and LLM evaluators. The first two may be explained by papers written for, or accepted at, a specific venue, and hence perhaps a higher quality bar. The latter, given its novelty, could be simply because of stronger claims.
When analysing the trends from Section 2.4, we observed decreases in citation rates between 2023 and 2024 for some criteria. This could have a simple explanation: newer papers have fewer citations. Other trends, such as the rise of the use of LLMs as evaluators and the increase on non-English evaluations, may be explained by our previous points on technological improvements, ease of access, and the previously-mentioned inflation of papers being produced.
4 Conclusion
We evaluated over 2,000 LLM-related scientific articles on criteria typical of what is considered good research (e.g., presence of statistical tests, reproducibility), and looked for arguments that have been at the centre of controversy and not present in reproducibility checklists for major conferences (e.g., claims of emergent behaviour, use of LLMs as evaluators).
We found various concerning trends. Yearly evaluations showed a decline on the presence of ethics disclaimers, open-sourcing, and statistical tests. Most papers using LLMs as evaluators claimed SOTA results. However, all papers claiming SOTA results relied less in statistical tests that those that did not. Claims that the models could reason were done with LLMs and not humans; contrasting with claims that they cannot not reason, which were typically carried out with human evaluation alone. All of this is evidence of overreliance and perhaps rushed research.
On the other hand, we observed a relatively steady amount of limitations sections, and an increasing number of evaluations in languages other than English. We attributed the former to hard requirements at conferences, and the latter to the advancement of this technology. In turn, this suggests that the combination of wide availability of powerful LLMs with enforcement of requirements does have a net positive effect on the literature, and on a more diverse and inclusive research community.
Other findings were a decrease on citations for papers claiming emergent behaviour, but a larger number of works claiming that the models can reason. We also found that there was a statistically significant difference in the likelihood for a paper to be cited if it had an ethics section present or used an LLM as an evaluator: we were unable to conclude if it was due to a higher quality bar for the paper, or because of the claims made.
Our findings underscore the need for more self-scrutiny and rigour by and from the field. This is no easy task. As shown here, LLM research is rapidly increasing in volume. Conferences are saturated with papers and scientists are encouraged to publish often. It is clear that critical reading beyond accepting prima facie the thesis behind a given work is crucial, now more than ever. Strong foundations on the field (e.g. statistics) and knowledge of the bleeding edge (e.g., findings supporting or contradicting the paper’s thesis) are a necessity to perform a judicious evaluation.
@@ 4 Conclusion
To close, Cremonesi and Jannach [8] mentioned, somewhat tongue-in-cheek, that in the context of recommender systems there was no reproducibility crisis. They posed that in a crisis researchers reflect upon and revise their methodologies. However, they argued, the field lay stagnant due to overfocusing on the same subjects without any introspection. It is the hope of this paper that the findings here encourage LLM researchers to keep pushing the field forward and living up to the fundamentals of a responsible scientific method: ethical, reproducible, systematic, and open to criticism.
5 Methods
5.1 Corpus
Our corpus is comprised of works that cited the peer-reviewed GPT-3 paper [36] (as opposed to the preprint) and the GPT-4 technical report [37]. We make the assumption that the majority of the LLM literature involves a reference to either of these papers.
We retrieved the top 1,000 papers sorted by citation numbers for both articles in Google Scholar [^4] https://scholar.google.com/ with Publish or Perish [42]; and the top 2,000 papers by citation number for GPT-3 in Scopus. [^5] https://www.scopus.com/ The disparity is due to Google Scholar indexing peer-reviewed works and preprints, and Scopus only indexing peer-reviewed articles. At the time of writing this paper, the GPT-4 technical report has yet to be peer-reviewed. All queries were ran for papers published or released up to 10 June 2024.
We then used the arXiv API [^6] https://info.arxiv.org/help/api/index.html to retrieve the full paper, and parsed either the HTML or the PDF source into text. Whenever these two operations failed we skipped the entry. In some instances, the title of the article and the results from the arXiv search query were mismatched. To mitigate this, our corpus cleanup and all experiments were done on the text, not the titles. The final, deduplicated, unlabelled corpus is $3,914$ texts.
5.2 Evaluation Criteria
We labelled our corpus based on a set of criteria (labels), categorised in four groups: Research Features, Structural Features, Arguments Made, and Indicators. They may be found in Table 1, with specific definitions–as prompts–in Appendix A.
Research Features, Structural Features, and Arguments Made are the core evaluation criteria used in the rest of our paper. Indicators is a filter for our corpus: we were only interested in papers that are research articles with an LLM as the subject of research.
5.3 Labelling
We labelled the data with GPT-4 Omni (version: gpt4-o-2024-05-13). To ensure reproducibility, we set the temperature to zero. We also set the maximum tokens to 256, and left other parameters as default. To improve accuracy we split in batches our labelling calls, totalling four different prompts [30]. All our calls were done through the Azure OpenAI API and the analysis done on a consumer-grade laptop.
The model was instructed to return binary labels ({y, n}) for all criteria, except for most Research Features labels, which also included a relevance label ({na}). To determine the evaluators used in a given paper we used the set {human, LLM, automatic, na}; and for the type of text, {book, article, opinion}.
We measured the model’s reliability by randomly sampling 100 papers per criterion, and manually labelling these. The model’s results were reliable to within an average 91.91 $\pm$ 1.22% with a 95% confidence interval. This number varies broadly across criteria, however. A full breakdown of reliability and analysis of performance is in Appendix B.
Prior to our experiments, we filtered the corpus by selecting all research articles with an LLM as the main subject of study. The final size of the data is $2,054$ papers.
@@ 6 Ethical Considerations
Open data is crucial for good research, but ethical and licencing considerations limit us from releasing the corpus with texts and personally-identifiable information. We release the code for our analysis under a permissive licence, and the annotated, anonymised data without texts. To avoid overloading the services we rate-limited our requests, and the crawling code will not be released.
@@ 7.1 Reliability of Automated Labelling
The community remains divided on the feasibility of using LLMs as evaluators. We argue that the reliability and conclusions drawn from using this technology vary with the problem and experimentation protocols used. We mitigate potential concerns by evaluating the performance of the model with statistical tests–and our analysis showed that GPT-4’s confidence bounds and accuracies indeed depended strongly on the criterion.
@@ 7.2 Corpus Representativeness
Our analysis is limited to works available on Google Scholar and Scopus citing the GPT-3 and GPT-4 papers. This might not represent the entire body of research on LLMs. As the literature and the technology evolves, we expect this assumption to hold less weight. However, as it stands, both papers are dominant in the literature. Due to time constraints we were unable to address an increasingly problematic issue in LLM research: synthetic data and use of possibly-contaminated benchmarks. We leave that exploration for future work.
@@ Acknowledgements
The author wishes to thank arXiv for use of its open access interoperability and A. Jangra and S. Visser for comments on an early version of this work.
@@ Appendix A Prompts
We split the labelling process in four to simplify the calls and because new criteria were added as the experimentation progressed. They can be found in Prompts 435 and  6. We used a single exemplar for each prompt. They were hand-picked from a paper that was not present in the corpus, and manually tuned for accuracy on a subset of the data ($n=10$) before labelling the full corpus. We requested the model to output a string during labelling. This string would be the rationale (for “na” labels), and the verbatim matching line of the paper otherwise. This technique has been shown to improve the model’s performance to out-of-distribution entries [43], and was helpful on analysing the performance of the model, which may be found in Appendix B.
Appendix B Labeller Reliability
Overall, the model performs well as a labeller, although certain criteria were certainly better-performing than others (e.g. open-sourcing versus SOTA claims). We partially attribute this to prompting. However, a closer inspection of the papers and the model’s rationale noted that the model tended to overlook content, sometimes verbatim, matching the criteria. The model had a tendency to make a liberal interpretation of the prompt: for example, for open sourcing, the model sometimes indicated that no open sourcing was performed because no LLMs were tested (which was not specified in the instructions; see Prompt 5). It also tended to frequently inject content about downloading the film “The Nun II” from Reddit.
Appendix C Extended Results
In this section we present the full results of our analysis, separated the same way as in Section 2: corpus composition results (Figures 5 and 6), composition over time (Figure 6), the relationship between citations and criteria (Table 8), and yearly trends on relationships between citations and criteria (Table 12). We also present selected plots for specific criteria.
Figure 5: Arguments Made (left) and Structural Features (right) for the papers claiming SOTA results. Overall we found a low prevalence of papers containing ethics disclaimers, error breakdowns, and negative results. Some of the arguments made (especially emergent behaviour) typically showed a much lower prevalence of the structural features considered to be good research.
Figure 6: Research quality for relevant papers claiming SOTA results. Although certain research features, such as non-English and dialect evaluations are very low, several standard research protocols, such as open-sourcing, versioning, and paper declarations are common in these papers.
@@ Appendix C Extended Results
For example,
<EXEMPLAR GOES HERE>
|begin paper|
<TEXT GOES HERE>
|end paper|
- emergence:
Table 3: Labelling prompt for the Indicators and parts of Arguments Made. Indicators had $99\%$ accuracy. LLM-as-a-subject and emergent behaviour had lower accuracy and looser confidence intervals ($89.0\pm 6.8$ and $83.0\pm 8.1$, respectively). Inspecting the output showed that the main cause of failure was GPT-4o returning (leaking) the exemplar for that criterion and ignored the input.
I am going to link a scientific paper. Tell me if the paper contains:
@@
- stat sig:
Table 4: Labelling prompt for part of Arguments Made and Structural Features. This prompt had good accuracy (between $93-100\%$) for all criteria with tight confidence intervals (±$\pm 0.0-5.5$).
@@
Answer everything with "y", "n", or "na", and add an explanation separated by a pipe. If you pick "y", return the verbatim first line matching.
For example,
<EXEMPLAR GOES HERE>
|begin paper|
<TEXT GOES HERE>
|end paper|
- versions:
Table 5: Labelling prompt for Research Features. The model had low performance on open-sourcing (74%), and acceptable (though low) accuracy in the other criteria. Examining GPT-4o’s reasoning for open-sourcing showed that it sometimes interpreted this label as only applicable if related to an LLM.
I am going to link a scientific paper. Tell me if the paper contains:
@@
- stat sig:
Table 6: Labelling prompt for the remaining Research Features. Non-English and dialects had good accuracy ($98$ and $100\%$). The type of evaluator had varying performances even though it was the same label: failures in human and LLM evaluators were mostly related to missing, rather than mislabelled, entries. Note that our analysis with human annotators has a smaller volume than LLM or automatic annotators.

### v3

@@ 4.2 What Changed Over Time?
Figure 3: Change over time for selected criteria. Between 2022 and 2024 there was an increase in the use of LLMs as evaluators; and declines on the presence of ethics disclaimers and open-sourcing. The use of human evaluators and limitation sections are steady.
For this analysis we worked with relevant papers claiming SOTA from 2021 onwards. Between 2023-2024 we observed declines in the absolute percentage of several criteria, such as the presence of ethics disclaimers, open-sourcing, claims of emergence, and statistical tests (Figure 3). There was an increase in the claims that these models can reason (15%); and a decrease in the claims that they cannot. Presence of limitation sections, error breakdowns, dialect evaluation, and relying on human evaluation remained steady ($\pm$1%). The use of LLM evaluators underwent an uptick (15%), but these papers generally have lower-than average proportions of research protocols. In terms of topics, papers have seen a doubling in subjects like multimodality and safety and security. Full results for topic changes are in Appendix C.3.
@@ 4.3 Do Papers With Certain Criteria Get More Citations?
We split our corpus in two (texts with and without a given criterion), and did a two-sample Kolmogorov-Smirnov test to determine the probability that both samples came from the same distribution. In this test, a high probability implies no significant difference between the samples. In this case we may then conclude that the presence of the criterion does not impact the number of citations, as they likely are drawn from the same distribution. Conversely, a low probability of being drawn from the same distribution allow us to conclude that the criterion is not related to the number of citations.
For our corpus we first calibrated the $p$-value to $<0.05$ for all criteria, which means that we expect to be wrong about our conclusions 5% of the time. If the test’s $p$-value is below our calibration threshold, we may reject the null hypothesis $H_{0}$ that the samples are related (i.e., that having one has an impact on the other). See Appendix C.4 for a detailed explanation of this test and full results.
Figure 4: Results of a Kolmogorov-Smirnov test on papers with (blue) and without (orange) ethics sections. The $p$-value for the test is p=$0.016$. This indicates that, if both samples were drawn from the same distribution, the probability that they are as far apart as observed is 1.6%. Since they were drawn from the same distribution, we may conclude that the presence of ethics sections does impact the number of citations received.
From our experiments, we rejected $H_{0}$ in the presence of ethics (Figure 4) and limitation sections; the use of LLM and automatic evaluators; open-sourcing; and claims of reasoning. This means that the presence of these on a paper had an impact on the citations it received. For other criteria (error breakdowns, evaluation of non-English languages, claims of emergence, and negative results) we were unable to reject $H_{0}$. Hence, their presence did not affect the citations received.
@@ 4.4 Are We Getting Better or Worse?
We analysed the relationship between citations and the presence of our criteria as yearly trends from 2021 onwards. We evaluated this relationship as the gap (absolute percentage difference) between citations of papers containing the criterion and these that do not, aggregated by year, and measured the change in this gap: positive changes imply that papers containing the criterion are cited more often. See Appendix C.5 for examples and the full results.
We observed an increase in the gap for ethics sections in 2022-2023 (+10%), but a noticeable drop in 2023-2024 (-50%). Other criteria had similar patterns, such as presence of statistical tests, limitation sections, and negative results. Note that 2024 gap drops are expected, given the recency of these papers. There were also upticks: LLMs as evaluators and non-English evaluations both had increases in citations. Unlike Section 4.3, recency bias is a concern in this experiment, which we discuss in the next section.
@@ 5 Discussion
5.1 Volume Analyses
Few papers claiming SOTA addressed ethical considerations, and this number is in decline. The proportions of statistical tests and open-sourcing in these papers are in line with critical literature for other CS fields. We also found open-sourcing and research protocols to be declining. Likewise, statistical tests are less often used than in LLM-related literature not claiming SOTA. Given the considerable volume of research contributed by the first half of 2024, these trends suggest rising rushed research and lack of rigour.
The use of LLMs as evaluators increased significantly, and most papers using them claimed SOTA results. This increase could be explained by technological developments (GPT-4, the standard of LLM-based evaluation, was released in 2023) and replication of experimental work. These works often coupled LLMs with other evaluator classes (e.g., humans), but eschewed measures of statistical significance or accounted for randomness in the calls. This also suggests rushed research, but at a lesser extent given the presence of other evaluators.
Steady criteria, namely limitation sections, may be explained by being a requirement of *ACL conferences Carpuat et al. (2024) since 2022: indeed, there is a $\sim$40increase of papers with this criterion in 2021-2022. The increased focuses on multimodality and safety indicate that this technology has matured beyond research; or, at least, that it is being widely adopted beyond NLP.
5.2 Claims
Claims of emergent behaviour were common–although in decline–but coupling them with statistical tests or error breakdowns were rare. On the other hand, claims that these models can reason were evaluated with LLMs and not humans; versus claims that they could not reason, typically done with human evaluation alone. Both suggest poor research practices: as per Section 2, metrics and evaluators are system-dependent, and a more robust approach would use more than one of both.
5.3 Citations and Criteria
We observed a statistically significant difference in the citations received by a paper when presenting ethics, limitations, and LLM evaluators. The first two may be explained by papers written for, or accepted at, a specific venue, and hence perhaps having higher quality. The latter, given its novelty, could be simply due to stronger claims. There was a statistically significant difference in the likelihood for a paper to be cited if it had an ethics section present or used an LLM as an evaluator. Due to likely recency bias, we were unable to conclude if it was because of a higher quality bar for the paper, the venue, or because of the claims made. It could be argued that ranking by citations induces a recency bias Kim et al. (2020); however, the (volume-based) results relying on citations as proxy do not change. Citations are a common metric to determine a paper’s success. There is also considerable skewness on their distribution, with 91% of the citations held by 25% of the corpus. Still, we caution on the interpretation of citation-based time-wise results.
5.4 Notes and Alternative Explanations
We observed decreases in citation rates in 2023-2024 for some criteria. This could have a simple explanation: newer papers have fewer citations. Hence we do not factor in these results in our conclusions. An uptick is not expected, however.
An alternative explanation for the decrease in ethics sections could be the multidisciplinary adoption of LLMs and the distinct ethical requirements for other fields. We consider matter-of-factly that ethics disclaimers in LLM-centred research have decreased, regardless of field. However, we do factor this observation into our recommendations, and discuss their relationships to venue requirements in Sections 6 and 9.
6 Conclusion and Takeaways
Our study is a critique of the methodology employed in LLM research. While statistically motivated, the centrepoints of our argument are related to taking stock of where we are and where are we going as a research community.
We noted a yearly decline on the proportion of papers having ethics disclaimers, open-sourcing, and statistical tests. Articles claiming SOTA relied more on LLMs as evaluators, and had fewer statistical tests, that the ones that did not. They also reported less research protocols, especially in 2024. We consider these evidence of mounting overreliance, lack of rigour, and/or rushed research, especially with respect to measurements.
That said, there were two major positive findings in our work. The first was a steady proportion of limitation sections, which we ascribed to requirements from conferences. The second was an increasing number of evaluations in non-English languages, and a rising interest on safety and security as a research subject. From both we conclude that a combination of available, powerful, LLMs; along with publication checklists does have a net positive effect on the literature, and on a more diverse and inclusive research community.
We also found that LLM-centred research is rapidly increasing in volume, and consequentially it could be that our assumption that most papers will cite GPT-4 or GPT-3 does not necessarily hold in the future. To determine this, a year after the data cutoff for our paper we attempted to perform a follow-up study. This could have determined any rates of change and recency biases induced by our measurements, along with the validation, or discrediting, of our core assumption. However, we were able to only validate the latter (Appendix D), because the tools we used in our work were no longer able to access Google Scholar. We argue that this is a concerning trend: the ability of surveys like ours to critically examine the field strongly rely heavily on open tools and publicly-available APIs. Losing that access could have ramifications on the ability of the field to perform self-examination.
Nonetheless, our work underscores the need for more self-scrutiny and rigour by and from the field. This is not easy given the overload of authors and peer-reviewers. It is also not feasible to wait months for the (purported) ‘next big thing’ to be peer reviewed if it is readily available as a preprint. While it could be said that critical reading is crucial, it would be naïve, however, to assume that this will be consistently done by all readers–e.g., laypeople, scientists not tasked with their peer review, etc. Our recommendations are designed to address this.
@@ 6 Conclusion and Takeaways
To close, Cremonesi and Jannach (2021) mentioned tongue-in-cheek that in the context of recommender systems there was no reproducibility crisis, as in a crisis researchers reflect upon and revise their methodologies. Instead, they stagnanted due to overfocusing on the same subjects without any introspection. It is our hope that our findings encourage LLM researchers to reflect on how to push the field forward, while also carrying out research that is ethical, systematic, and open to criticism.
7 Recommendations
Based on our observations from the previous section, we extend three recommendations. These call for specific features in papers that should be scrutinised during peer review. This is because, when researchers know that either the venue will enforce some features (e.g., limitation sections), or, that reviewers will ask for them (e.g., releasing artifacts), they will usually add them to the pre-review preprints. While this is not common outside of CS, adding these features within NLP will encourage researchers to reflect upon their work and allow other readers to readily comprehend the scope of the research. Hence, we summarise our recommendations to the scientific community in three areas: impact analysis, measurement rigour, and transparency.
Impact Analysis: Venues should continue (or begin) to enforce (short) sections that allow for easier critical evaluation of the work, and encourage self-reflexion and thoughtful research by the authors. The two sections are (1) limitations/scope, and (2) considerations on the broader impact of the work. The first must disclose the experimentation’s boundaries and areas of improvement/future work. The second is not an ethics disclaimer, but still allows readers to understand the implications of the research. Peer-review feedback after reviews should be added in: as third parties without any stakes in the work, they are best positioned to provide informed critiques that other readers may miss, even if available elsewhere (e.g., OpenResearch). Neither section should count towards the page limit.
Measurement Rigour: Reviewer checklists should explicitly account for (but not require) statistical tests, [^7] Their definition must also be included in the work. number and type of metrics and languages evaluated, and classes of evaluators used. We do not call for their enforcement as that these are context-dependent. That said, since LLM work is primarily empirical and comparison-based, they are likely going to be required often.
The idea is that measurement rigour should explicitly be part of an even evaluation of a paper’s merits. It could also allay some of the concerns from Section 2 on system-dependence and reliability, and propose areas for further work. This does not mean that papers using LLM evaluators alone should be discounted outright, but their methodologies should be scrutinised closely. This is especially important for works using a single prompt or LLM evaluator, such as this one. Without a meta-evaluation, the results should be deemed unreliable.
Transparency in LLM research is tricky, but not unattainable. Terminology (e.g., AGI, emergence, reasoning) must be formally and carefully defined if evaluated in the work. Declarations of prompts, call parameters, and versioning must be enforced. Reviewers should be encouraged to seek an error analysis section focused on the LLMs’ responses: LLMs are notoriously unreliable, and they are better understood when analysing their responses qualitatively. On open-sourcing, it is worth noting that a model’s weights being open-sourced does not constitute transparency, if neither the code nor the data were released. Unlike earlier recommendations, this is a matter of semantics and may be caught during peer review.
At a more meta level–and not strictly related to peer-review–our finding that public APIs could not access certain sites signifies a loss of transparency that could have repercussions on the ability of the field to perform self-examination. We recommend that these APIs should have their access re-enabled.
@@ 8 Ethical Considerations
Open data is crucial for good research, but ethical and licencing considerations limit us from releasing the corpus with texts and personally-identifiable information. We release the code for our analysis under a permissive licence (MIT), and the annotated, anonymised data without texts. To avoid overloading the services we rate-limited our requests, in compliance with their terms of use; and the crawling code will not be released.
@@ 9.1 Reliability of Automated Labelling
The community remains divided on the feasibility of using LLMs as evaluators. We argue that the reliability and conclusions drawn from using this technology vary with the problem and experimentation protocols used. We mitigate potential concerns by evaluating the performance of the model with statistical tests: our analysis showed that GPT-4’s confidence bounds and accuracies were reliable.
@@ 9.2 Corpus Representativeness
Our analysis is limited to works available on Google Scholar and Scopus citing the GPT-3 and GPT-4 papers. This might not represent the entire body of research on LLMs. As the literature and the technology evolves, we expect this assumption to hold less weight. However, as it stands, a year after the cutoff data for our paper, both papers are still very dominant in the literature. See Appendix D for a follow-up study evaluating our assumption.
This also in turn overlooks a potential issue with citation-based ranking: well-known authors, venues, and institutions could have more citations just by virtue of being known. They could also evolve over time, inducing sampling biases Kim et al. (2020). This is, however, a common shortcoming to papers like ours (c.f., Pramanick et al. 2023), and, as argued, not a detractor from the core points of our work.
9.3 Venues in Scope
One of the main findings of this work was that the requirements from some venues (e.g., ACL) around mandatory sections were successful at maintaining their presence stable. However, a careful study could distinguish between venues to ablate out correlations. Nonetheless, the experimental setup, relying on open APIs, means that a large volume of papers evaluated may not necessarily be accepted, or even submitted, to these venues. This makes said experimentation tricky and likely the subject of future work.
9.4 Criteria Established
Due to time constraints we were unable to address an increasingly problematic issue in LLM research: synthetic data and use of possibly-contaminated benchmarks. We leave that exploration for future work.
@@ Acknowledgements
The author wishes to thank arXiv for use of its open access interoperability; A. Jangra and S. Visser for comments on an early version of this work; and the anonymous reviewers, whose informed critiques were fundamental to improving this work and were incorporated into the final version.
@@ Appendix A Prompts
We split the labelling process to simplify the calls and because new criteria were added as the experimentation progressed. They can be found in Prompts 1 and 2 (criteria) and Prompt 3 (topic analysis). We used a single exemplar for each prompt. They were hand-picked from a paper that was not present in the corpus, and manually tuned for accuracy on a subset of the data ($n=10$) before labelling the full corpus. We requested the model to output a string during labelling. This string would be the rationale (for ‘na’ labels), and the verbatim matching line of the paper otherwise. This technique has been shown to improve the model’s performance to out-of-distribution entries Brahman et al. (2021), and was helpful on analysing the performance of the model, which may be found in Appendix B.
@@
Table 1: Labelling prompt for the Indicators and parts of Arguments Made (in blue) and Structural Features and the remaining Arguments Made (in red). Other areas are shared between both prompts; but blue lines do not appear in red, and viceversa. Indicators had $99\%$ accuracy. LLM-as-a-subject and emergent behaviour had lower accuracy and looser confidence intervals ($89.0\pm 6.8$ and $83.0\pm 8.1$, respectively). Inspecting the output showed that the main cause of failure was GPT-4o returning (leaking) the exemplar for that criterion and ignored the input. The red prompt had good accuracy (between $93-100\%$) with tight confidence intervals (±$\pm 0.0-5.5$).
@@
Answer everything with "y" or "n", or "na", and add an explanation separated by a pipe. If you pick "y", return the verbatim first line matching.
@@
Table 2: Labelling prompt for the first (blue) and second (red) parts of Research Features. The model had low performance on open-sourcing (74%), and acceptable (though low) accuracy in the other criteria. Non-English and dialects had good accuracy ($98$ and $100\%$). Examining GPT-4o’s reasoning for open-sourcing showed that it sometimes interpreted this label as only applicable if related to an LLM. The type of evaluator had varying performances even though it was the same label: failures in human and LLM evaluators were mostly related to missing, rather than mislabelled, entries.
I am going to link a scientific paper. Tell me what is the primary and secondary subject of the paper, in at most three words each.
Return the extract of the paper, verbatim, that correspond to why you picked that subject.
For example,
<EXEMPLAR GOES HERE>
|begin paper|
<TEXT GOES HERE>
|end paper|
- primary:
Table 3: Topic clustering prompt. The output distribution was very wide, with about 1,000 lexically unique entries. Manual verification was needed for most topics to reduce and classify it to our 10 primary topics.
Appendix B Labeller Reliability
Overall, the model (prompt) performs well as a labeller, although certain criteria were certainly better-performing than others (e.g. open-sourcing versus SOTA claims). We partially attribute this to prompting. However, a closer inspection of the papers and the model’s rationale noted that the model tended to overlook content, sometimes verbatim, matching the criteria. The model had a tendency to make a liberal interpretation of the prompt: for example, for open sourcing, the model sometimes indicated that no open sourcing was performed because no LLMs were tested (which was not specified in the instructions; see Prompt 2). It also tended to frequently inject content about downloading the film ‘The Nun II’ from Reddit.